QUESTION 3:

How much does Savannah need to save at the end of each of the next twenty-five years (with the first deposit made at the end of year 1) to have an income of $40,000 per year for thirty years starting at end of year twenty-six. The appropriate interest rate is 6.0% p.a.

QUESTION 4:

To buy an engagement ring for his girlfriend, Martin is considering a loan of $10,000 from Tiffney. He will pay 36 monthly installments of $550 each. What is the effective annual rate of interest (EAR) that Tiffney is charging?
Answer #1

1.

x/0.06*(1-1/1.06^25)=40000/1.06^26*(1-1/1.06^30)/(1-1/1.06)=

x=10035.508

2.

=(1+RATE(36,-550,10000,0)*12/12)^12-1

=65.47%